OpenStudy (anonymous):

Write the equation of the line with slope = - 3 and containing the point (- 7, 6).

OpenStudy (anonymous):

y - 6 = -3(x+7)

OpenStudy (anonymous):

ok that isn't one of the answers, so I will choose non of the above!

OpenStudy (anonymous):

use x-a=m(y-b) where (a,b) is the point and m is slope ..here (-7,6) and -3..put these value in the eqn and get the ans

OpenStudy (anonymous):

Thank you both!